40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 19 17:11:52 2014, steve, Update, PEM, it is warmer today warmertoday.png
    Reply  Mon Sep 22 10:18:58 2014, steve, Update, PEM, lab temperatures summerheat.png
       Reply  Tue Nov 4 10:40:46 2014, steve, Update, PEM, lab temperatures and particle counts PEM100d.png
          Reply  Mon Dec 8 09:03:49 2014, steve, Update, PEM, air cond maintenance tomorrow morning 
             Reply  Wed Dec 10 14:26:13 2014, steve, Update, PEM, air cond maintenance and particle plot 05.png
                Reply  Fri Sep 16 14:15:28 2016, steve, Update, PEM, partical counts 80days.png
                   Reply  Fri Dec 2 10:57:08 2016, steve, Update, PEM, door seal replaced 1_year_PEM.png
                   Reply  Fri Feb 17 10:56:12 2017, steve, Update, PEM, particle counter moved into PSL enclousure from_here.jpgto_here.jpgPSL_particles.png
                      Reply  Tue Feb 21 10:08:35 2017, steve, Update, PEM, noisy morning outside_activity.png
                      Reply  Wed Feb 22 08:54:17 2017, steve, Update, PEM, PSL enclousure particle count dusty__PSL.png
                         Reply  Tue Feb 28 08:04:48 2017, steve, Update, PEM, ETMX damping recovered ETMX.png
                            Reply  Tue Feb 28 18:25:22 2017, rana, Update, PEM, ETMX damping recovered 
                               Reply  Thu Mar 16 08:22:16 2017, Steve, Update, SUS, ETMX damping  laser_power_glitch.png
                            Reply  Mon Aug 28 10:51:21 2017, steve, Update, SUS, ETMX damping recovered ETMX_restored.png
                      Reply  Mon Mar 20 17:12:08 2017, steve, Update, PEM, particle counter inside of PSL enclousure enclousure_partical_count.png
                         Reply  Mon Mar 27 15:36:27 2017, Steve, Update, PEM, particle counter inside of PSL enclousure HEPA_flow_effects.png
Message ID: 12908     Entry time: Mon Mar 27 15:36:27 2017     In reply to: 12895
Author: Steve 
Type: Update 
Category: PEM 
Subject: particle counter inside of PSL enclousure 

 

Quote:

The logging script is multiplying by 100 instead of 10 !

Quote:

The MET#1 particle counter was moved from CES wall at ITMX to PSL enclousure south west corner at 11am.

The HEPA filter speed at the Variac was turned down to 20V from 40

This counter pumps air for 1 minute in every 20 minutes. Soft foam in bags used to minimize this shaking as it is clamped.

 

 

Some one turned up the PSL HEPA rotation voltage from 20 V to 33 ( 120V Variac ) and  X-arm AC set temp lowered to 68F

Effects at 12" height from optical table :

                                                            1.0, 0.5 & 0.3 micron particle count  goes to zero and temperature fluctuaion increased

Rotation speed voltage was set to 30V

                                                            

Attachment 1: HEPA_flow_effects.png  35 kB  Uploaded Mon Mar 27 17:15:17 2017  | Hide | Hide all
HEPA_flow_effects.png
ELOG V3.1.3-